OpenStudy (anonymous):

Algebra 2 Help??? simplify: (8/(c^2-4))+(2/(c^2-5c+6))

OpenStudy (anonymous):

thats gotta equal something

HanAkoSolo (jamierox4ev3r):

lol what kind of answer is that?!

OpenStudy (anonymous):

\[\frac{ 8 }{c^2-4 }+\frac{ 2 }{c^2 -5c+6 }=\frac{ 8 }{(c+2) (c-2) }+\frac{ 2 }{ (c-2)(c-3) }\] can you continue?

OpenStudy (anonymous):

\[=\frac{ 8(c-3) }{(c +2)(c-2) (c-3)}+\frac{ 2(c+2) }{(c+2)c-2)(c-3) }\]

OpenStudy (anonymous):

the leftover is yours